# Entropy of self-avoiding branching polymers: mean field theory and Monte Carlo simulations

### v1

@@ Abstract
We study the statistics of branching polymers with excluded-volume interactions, by modeling them as single self-avoiding trees on a generic regular periodic lattice with coordination number $q$. Each lattice site can be occupied at most by one tree node, and the fraction of occupied sites can vary from dilute to dense conditions. By adopting the statistics of directed trees as a proxy for that of undirected trees without internal loops and by an exact mapping of the model into a field theory, we compute the entropy and the mean number of branch-nodes within a mean field approximation and in the thermodynamic limit. In particular, we find that the mean number of branch-nodes is independent of both the lattice details and the lattice occupation, depending only on the associated chemical potential. Monte Carlo simulations in $d=2,3,4$ provide evidence of the remarkable accuracy of the mean field theory, more accurate for higher dimensions.
@@
Approximate theories were proposed in the past to study interacting branching polymers Lubensky and Isaacson (1979, 1981); Nemirovsky, Dudowicz, and Freed (1992); Gujrati (1995); Banchio and Serra (1995); De Los Rios, Lise, and Pelizzola (2001); Zilman and Safran (2002); Wagner, Erdemci-Tandogan, and Zandi (2015). In most cases Lubensky and Isaacson (1979, 1981); Nemirovsky, Dudowicz, and Freed (1992); Zilman and Safran (2002); Wagner, Erdemci-Tandogan, and Zandi (2015) they consist in mean field (hereafter, MF) approximations of lattice field theories constructed upon the “$n\to 0$”-limit of the spin $O(n)$-model of a magnet, following the original idea for linear chains proposed first by Daoud et al. Daoud et al. (1975) and formalized later by de Gennes de Gennes (1979). Yet, these classical works study systems where the presence of loops cannot be truly avoided, and the number and size of the polymer molecules cannot be controlled. Another possible approximation consists in embedding the polymer system into a Bethe lattice Gujrati (1995); Banchio and Serra (1995). In this framework, the absence of loops in the polymer structure is ensured by the loopless nature of the underlying lattice itself. However, the constraint that a polymer in real space (or in a conventional periodic lattice) must avoid loops introduces a nontrivial correlation between nodes, which is neglected in the Bethe lattice representation. An alternative attempt to construct a theory for single-chain systems avoiding loops on a standard periodic lattice was proposed in De Los Rios, Lise, and Pelizzola (2001): here, however, the corresponding MF free energy requires numerical minimization and the amount of branch-nodes in the system cannot be controlled, keeping into the same ensemble configurations of linear chains and highly-branched structures.
Motivated by these considerations, in this Communication we introduce a model for branching polymers with explicit excluded-volume interactions which solves all these issues, namely: (i) loops are rigorously excluded, thus providing a general framework for naturally occurring systems like those mentioned at the beginning; (ii) the approach is genuinely single-chain; (iii) control is maintained on both the amount of branch-nodes (and, so, explore the whole range from the poorly- to the highly-branched regime) and monomer concentration. We then proceed via an exact mapping of the partition function on a field theory which can be solved within a mean field approximation. This solution includes many other known results as limit cases and allows us to derive a transparent analytical formula for the mean number of branch-nodes of the trees which is in semi-quantitative agreement with known predictions from exact-enumeration techniques and new Monte Carlo computer simulations of self-avoiding trees on the hypercubic lattice in $d=2,3,4$ spatial dimensions.
@@
The model – We consider a regular lattice with coordination number $q$, total number of sites $V$, and where periodic boundary conditions are assumed. A single randomly branching polymer with excluded-volume interactions is modeled as a self-avoiding tree (SAT) made of $N\leq V$ monomers (nodes), or $N-1$ bonds, on such lattice. A site of the lattice can be occupied by at most one tree node and the bond between two nodes (equal to one lattice step) corresponds to our length unit (see Fig. 1(a) for a simple SAT conformation on the square lattice). In the absence of loops, denoting with $f_{i}$ the number of bonds emanating from node $i$ (the so called functionality) and with $N_{f}$ the total number of nodes of functionality $f$, we have that $N=\sum_{f=1}^{q}N_{f}$ while the $f_{i}$’s sum to twice the total number of bonds, i.e. $\sum_{i=1}^{N}f_{i}=\sum_{f=1}^{q}f\,N_{f}=2(N-1)$. Then, from these relations one obtains:
@@
where the upper limit $f=q$ in the sums follows from the self-avoidance condition. [^1] Of course, for ideal trees with no excluded volume nothing prevents $f$ to be $>q$. Notice also that the obvious relation $N=\sum_{f=1}^{q}N_{f}$ follows immediately from Eqs. (1) and (2).
SAT entropy, MF approximation – Our task consists in computing $\Omega_{N}(\left\{N_{f}\right\})$, the total number of SAT’s of $N$ nodes and $\left\{N_{f}\right\}\equiv(N_{3},...,N_{q})$ branch-nodes. To this purpose, we consider the related problem of the so called directed-rooted SAT’s, namely trees where every bond has an assigned orientation such that there exists a unique directed path from a “privileged” node (called root), see panel (b) in Fig. 1. With the root taken on the (arbitrarily chosen) lattice site position ${\mathbf{x}}_{0}$, the corresponding total number of directed-rooted SAT’s, $\Omega_{N}^{{\mathbf{x}}_{0}}(\left\{N_{f}\right\})$, is related to $\Omega_{N}(\left\{N_{f}\right\})$ as shown by the following argument. With reference to Fig. 1, the degeneracy due to translational and rotational symmetry is $=qV$ and $=qN$ for undirected and directed trees, respectively. It follows then, that the total number of distinct tree conformations in the two cases is $=\Omega_{N}/(qV)$ and $=\Omega_{N}^{{\mathbf{x}}_{0}}/(qN)$, respectively. As these two last quantities must be identical [^2] Notice that the assumption of periodic boundary conditions guarantees the equality irrespective of the choice of ${\mathbf{x}}_{0}$., by introducing the Boltzmann constant $k_{B}$ we get the relation:
@@
between the entropies per lattice site, $s\equiv k_{B}\log(\Omega_{N})/V$ and $s^{{\mathbf{x}}_{0}}\equiv k_{B}\log(\Omega_{N}^{{\mathbf{x}}_{0}})/V$, for undirected and directed SAT’s respectively. In the thermodynamic limit $V\gg 1$ we can neglect terms of order $1/V$, and Eq. (3) implies $s\simeq s^{{\mathbf{x}}_{0}}$. This observation is now crucial because, contrary to $\Omega_{N}$, $\Omega_{N}^{{\mathbf{x}}_{0}}$ can be mapped exactly on a field theory, [^3] Essentially, the reason why we get “easily” $\Omega_{N}^{{\mathbf{x}}_{0}}$ but not $\Omega_{N}$ is because by assigning a direction to each bond, looped conformations have two bonds pointing to the same site: in this way (see Sec. S1 in SM) their statistical weight is $=0$ and are so naturally removed from the count. At the same time, by our rules it is not too difficult to demonstrate (Sec. S2 in SM) that an alternative field theory like the one proposed by Wagner et al. Wagner, Erdemci-Tandogan, and Zandi (2015) does, in fact, count loops. for which we can adopt a scheme similar to that recently introduced by us Marcato et al. (2023, 2024) to model interacting linear polymer solutions and largely based on the well-known de Gennes’ mapping of the “$n\to 0$”-limit of the spin $n$-vector model de Gennes (1979) (the complete derivation of $\Omega_{N}^{{\mathbf{x}}_{0}}$ is given in Sec. S1 in Supplementary Material (SM)). The main idea consists in writing the grand canonical partition function of the system $\Xi$ in the following field-theoretic form,
@@
where $\bm{\varphi}$ and $\bar{\bm{\varphi}}$ are $n$-component field vectors and ${\mathcal{A}}({\bm{\varphi}},\bar{\bm{\varphi}})$ is an effective action (see Eq. (S17) in SM). While the final expression remains mathematically challenging, it is particularly amenable to an expansion around the saddle-point Orland, Itzykson, and de Dominicis (1985); Doniach, Garel, and Orland (1996); Higuchi (1998); Marcato et al. (2023, 2024), which is obtained by minimizing the effective action. The first term of this expansion corresponds to our MF estimate. Finally, a standard Legendre transform allows us to switch back to the canonical ensemble and get finally $\Omega_{N}^{{\mathbf{x}}_{0}}$. By neglecting all terms up to ${\mathcal{O}}(V^{-1})$, the MF expression for $s=s_{\rho}(\left\{\phi_{f}\right\})$ (Eq. (3)) for finite node density $\rho\equiv N/V$ and as a function of branch-node fractions $\left\{\phi_{f}\right\}\equiv(\phi_{3},...,\phi_{q})$ with $\phi_{f}\equiv N_{f}/N$ is given by
@@
with $e=2.718...$ being the Euler number, ()!(...)!( … ) ! the factorial symbol and with the constraints $\phi_{1}=\sum_{f=3}^{q}(f-2)\phi_{f}$ and $\phi_{2}=1-\sum_{f=3}^{q}(f-1)\phi_{f}$ following, respectively, from Eqs. (1) and (2).
@@
i.e. the entropy per node $s_{\rho}/\rho$ is constant. Interestingly, the argument of the logarithm plays now the role of an effective coordination number Lipson and Whittington (1983) which depends on the set of branching probabilities $\{\phi_{f}\}$. In particular, this effective coordination number can take smaller values with respect to the one for linear chains (which, within our MF approximation, is $=q$, i.e. the same as for random walks on the lattice): for instance, by setting $\phi_{3\leq f<q}=0$ and $\phi_{q}\neq 0$, for $q\geq 6$, there are values of $\phi_{q}$ giving an effective coordination number $<q$. Third, of even greater interest is the opposite limit of spanning trees, $\rho=1$, namely at full lattice occupation. In particular, through the Legendre transform of Eq. (5) (see Sec. S3 in SM) it is possible to enumerate the total number of spanning trees on the lattice with unrestricted $N_{f}$. Our MF estimate reproduces the known Wu (1977); Shrock and Wu (2000) exact entropies of spanning trees for different lattices with good accuracy (deviations are no larger than $\simeq 20\%$, see Table S1 in SM). While our approximation is less precise than the Bethe approximation of De Los Rios et al. De Los Rios, Lise, and Pelizzola (2001) (see Table S1 in SM), our method is conceptually superior as it allows us to seize control of the mean number of branch-nodes, a feature absent in De Los Rios, Lise, and Pelizzola (2001).
@@
Figure 2: Distribution of equilibrium densities $\langle\phi_{f}\rangle$ for nodes of functionality $f$ in spanning trees on the triangular lattice ($q=6$). Orange columns correspond to our MF result for an infinite lattice, whereas blue columns refer to the results obtained through the numerical procedure of Pozrikidis Pozrikidis (2016) for a periodic lattice with a total number of $256$ sites.
@@
Figure 3: Asymptotic mean fraction of branch-nodes, $\langle\phi_{3}\rangle$, as a function of the branch chemical potential $\beta\mu_{3}$, comparison between MF prediction Eq. (9) (dashed line) and MC numerical simulations (symbols, corresponding to the values for the largest $N$’s reported in Fig. S3 in SM) for different node densities $\rho$ (see colorcode/symbol in the legend) and in spatial dimensions $d=2,3,4$. (Insets) Corresponding ratios between MC results and the MF prediction.
@@
where $\beta=1/(k_{B}T)$. Inverting Eq. (8) for given $\{\mu_{f}\}$, the unknown mean values $\{\langle\phi_{f}\rangle\}$ obey a system of $q-2$ equations which, in general, needs to be solved numerically. However, a general feature easily deduced from Eqs. (5) and (8) is that the values of $\{\langle\phi_{f}\rangle\}$ do not depend on $\rho$. We study first the case of spanning trees (i.e. $\rho=1$) with unrestricted functionality on the triangular lattice (i.e. $q=6$ and $\beta\mu_{f}=0$ for $3\leq f\leq 6$) for which accurate numerical predictions are available Pozrikidis (2016). As shown in Fig. 2, our MF theory agrees very well with the numerical results. Then we consider the situation where $\beta\mu_{f}\neq 0$ and assume that branch-nodes are at most $3$-functional regardless of lattice coordination $q$. With that, Eq. (8) reduces to the following simple expression for $\langle\phi_{3}\rangle$ as a function of the chemical potential $\mu_{3}$:
@@
[…]
@@
Conclusions – In this work, we have presented a field-theoretic method to study the thermodynamics of single, self-avoiding trees on a generic, regular lattice. We have obtained a compact mean field expression (Eq. (5)) for the SAT entropy that is, at the same time, significantly simpler and more general than others obtained in seemingly MF contexts Lubensky and Isaacson (1979, 1981); De Los Rios, Lise, and Pelizzola (2001); Zilman and Safran (2002); Wagner, Erdemci-Tandogan, and Zandi (2015), and whose accuracy (see Table S1 in SM) can in principle be improved by the systematic inclusion of higher order terms beyond the saddle-point approximation utilized here. Notably, our result provides a simple, analytical formula that is able to interpolate between dilute trees ($\rho\ll 1$) and spanning trees ($\rho=1$), as well as between linear and highly branched structures. The results for all these different scenarios have been compared with exact results from the literature, highlighting the good accuracy of our MF approximation. We have also provided predictions for the branching probability (Eq. (9)) that are in good agreement with computational results from Monte Carlo simulations. In this regard, a new simple, yet effective algorithm for the simulation of spanning trees has also been presented. Finally, with little additional effort the model can be extended to include the contribution of nearest-neighbor attractive interactions between tree nodes to model the situation of branched polymers in poor-solvent conditions Rubinstein and Colby (2003). In particular (see Sec. S5 in SM), poor-solvent interactions do not affect the mean values of branch-node fractions found for simple excluded-volume conditions; at the same time, the expected globule-coil transition as a function of temperature is of the same kind of the one for simple linear (i.e. unbranched) polymers in agreement with works Nemirovsky, Dudowicz, and Freed (1992); De Los Rios, Lise, and Pelizzola (2001).
@@
Entropy of self-avoiding branching polymers: mean field theory and Monte Carlo simulations
@@
S1 Field-theoretic partition function and mean field approximation
In this Section, we illustrate in detail the derivation of the mean field expression of the entropy (Eq. (5) in main text) for self-avoiding trees (SAT’s).
S1.1 Spin $n$-vector model for SAT’s and $n\to 0$ limit
S1.1.1 Formulation of the problem
We consider a regular lattice with coordination number (equal to the number of nearest neighbors per lattice site) $=q$, total number of sites $=V$ and with periodic boundary conditions; lengths are expressed in units of the lattice step length. As explained in main text, the goal is to estimate $\Omega_{N}^{{\mathbf{x}}_{0}}(\left\{N_{f}\right\})$, the total number of rooted-directed trees with root on the lattice site ${\mathbf{x}}_{0}$, $N$ nodes and $\left\{N_{f}\right\}\equiv(N_{3},...,N_{q})$ branch-nodes of functionality $f$.

### v2

# Entropy of self-avoiding branching polymers: Mean-field theory and Monte Carlo simulations

@@ Abstract
We study the statistics of branching polymers with excluded-volume interactions, by modeling them as single self-avoiding trees on a generic regular periodic lattice with coordination number $q$. Each lattice site can be occupied at most by one tree node, and the fraction of occupied sites can vary from dilute to dense conditions. By adopting the statistics of rooted-directed trees as a proxy for that of undirected trees without internal loops and by an exact mapping of the model into a field theory, we compute the entropy and the mean number of branch-nodes within a mean-field approximation and in the thermodynamic limit. In particular, we find that the mean number of branch-nodes is independent of both the lattice details and the lattice occupation, depending only on the associated chemical potential. Monte Carlo simulations in $d=2,3,4$ provide evidence of the remarkable accuracy of the mean-field theory, more accurate for higher dimensions.
@@
Approximate theories were proposed in the past to study interacting branching polymers Lubensky and Isaacson (1979, 1981); Nemirovsky, Dudowicz, and Freed (1992); Gujrati (1995); Banchio and Serra (1995); De Los Rios, Lise, and Pelizzola (2001); Zilman and Safran (2002); Wagner, Erdemci-Tandogan, and Zandi (2015). In most cases Lubensky and Isaacson (1979, 1981); Nemirovsky, Dudowicz, and Freed (1992); Zilman and Safran (2002); Wagner, Erdemci-Tandogan, and Zandi (2015) they consist in mean-field (hereafter, MF) approximations of lattice field theories constructed upon the $n\to 0$”-limit of the spin $O(n)$-model of a magnet, following the original idea for linear chains proposed first by Daoud et al. Daoud et al. (1975) and formalized later by de Gennes de Gennes (1979). Yet, these classical works study systems where the presence of loops cannot be truly avoided, and the number and size of the polymer molecules cannot be controlled. Another possible approximation consists in embedding the polymer system into a Bethe lattice Gujrati (1995); Banchio and Serra (1995). In this framework, the absence of loops in the polymer structure is ensured by the loopless nature of the underlying lattice itself. However, the constraint that a polymer in real space (or in a conventional periodic lattice) must avoid loops introduces a nontrivial correlation between nodes, which is neglected in the Bethe lattice representation. An alternative attempt to construct a theory for single-chain systems avoiding loops on a standard periodic lattice was proposed in De Los Rios, Lise, and Pelizzola (2001): here, however, the corresponding MF free energy requires numerical minimization and the amount of branch-nodes in the system cannot be controlled, keeping into the same ensemble configurations of linear chains and highly-branched structures.
Motivated by these considerations, in this Communication we introduce a model for branching polymers with explicit excluded-volume interactions which solves all these issues, namely: (i) loops are rigorously excluded, thus providing a general framework for naturally occurring systems like those mentioned at the beginning; (ii) the approach is genuinely single-chain; (iii) control is maintained on both the amount of branch-nodes (and, so, explore the whole range from the poorly- to the highly-branched regime) and monomer concentration. We then proceed via an exact mapping of the partition function on a field theory which can be solved within a mean-field approximation. This solution includes many other known results as limit cases and allows us to derive a transparent analytical formula for the mean number of branch-nodes of the trees which is in semi-quantitative agreement with known predictions from exact-enumeration techniques and new Monte Carlo computer simulations of self-avoiding trees on the hypercubic lattice in $d=2,3,4$ spatial dimensions.
@@
The model – We consider a regular lattice with coordination number $q$, total number of sites $V$, and where periodic boundary conditions are assumed. A single randomly branching polymer with excluded-volume interactions is modeled as a self-avoiding tree (SAT) made of $N\leq V$ monomers (nodes), or N1N-1 bonds, on such lattice. A site of the lattice can be occupied by at most one tree node and the bond between two nodes (equal to one lattice step) corresponds to our length unit (see Fig. 1(a) for a simple SAT conformation on the square lattice). In the absence of loops, denoting with $f_{i}$ the number of bonds emanating from node $i$ (the so called functionality) and with $N_{f}$ the total number of nodes of functionality $f$, we have that $N=\sum_{f=1}^{q}N_{f}$ while the $f_{i}$’s sum to twice the total number of bonds, i.e. $\sum_{i=1}^{N}f_{i}=\sum_{f=1}^{q}f\,N_{f}=2(N-1)$. Then, from these relations one obtains:
@@
where the upper limit $f=q$ in the sums follows from the self-avoidance condition. [^1] Of course, for ideal trees with no excluded volume nothing prevents $f$ to be >q>q. Notice also that the obvious relation $N=\sum_{f=1}^{q}N_{f}$ follows immediately from Eqs. (1) and (2).
SAT entropy, MF approximation – Our task consists in computing $\Omega_{N}(\left\{N_{f}\right\})$, the total number of SAT’s of $N$ nodes and $\left\{N_{f}\right\}\equiv(N_{3},...,N_{q})$ branch-nodes. To this purpose, we consider the related problem of the so called rooted-directed SAT’s, namely trees where every bond has an assigned orientation such that there exists a unique directed path from a “privileged” node (called root), whose spatial position is taken at the (arbitrarily chosen) fixed lattice site position ${\mathbf{x}}_{0}$ (see panel (b) in Fig. 1). Then, the total number of rooted-directed SAT’s, $\Omega_{N}^{{\mathbf{x}}_{0}}(\left\{N_{f}\right\})$, is related to $\Omega_{N}(\left\{N_{f}\right\})$ as shown by the following argument. As a direct consequence of (i) assuming periodic boundary conditions and (ii) fixing the root position ${\mathbf{x}}_{0}$, each distinct tree conformation has a degeneracy due to translations equal to $V$ in the case of undirected configurations, and equal to $N$ in the case of rooted-directed configurations. In fact, for a rooted-directed SAT configuration there must always be a monomer (i.e. the root) in position $\mathbb{x}_{0}$, thus limiting the possible translations. The degeneracy due to rotations is instead the same for both undirected and rooted-directed configurations (details, in particular Fig. S1, in Sec. S1 in Supplementary Material (SM)). Hence the ratio $\Omega_{N}(\left\{N_{f}\right\})/\Omega_{N}^{{\mathbf{x}}_{0}}(\left\{N_{f}\right\})=V/N$. Then, by introducing the Boltzmann constant $k_{B}$ we get the relation:
@@
between the entropies per lattice site, $s\equiv k_{B}\log(\Omega_{N})/V$ and $s^{{\mathbf{x}}_{0}}\equiv k_{B}\log(\Omega_{N}^{{\mathbf{x}}_{0}})/V$, for undirected and rooted-directed SAT’s respectively. In the thermodynamic limit $V\gg 1$ we can neglect terms of order $1/V$, and Eq. (3) implies $s\simeq s^{{\mathbf{x}}_{0}}$. This observation is now crucial because, contrary to $\Omega_{N}$, $\Omega_{N}^{{\mathbf{x}}_{0}}$ can be mapped exactly on a field theory, [^2] Essentially, the reason why we get “easily” $\Omega_{N}^{{\mathbf{x}}_{0}}$ but not $\Omega_{N}$ is because by assigning a direction to each bond, looped conformations have two bonds pointing to the same site: in this way (see Sec. S2 in SM) their statistical weight is =0=0 and are so naturally removed from the count. At the same time, by our rules it is not too difficult to demonstrate (Sec. S3 in SM) that an alternative field theory like the one proposed by Wagner et al. Wagner, Erdemci-Tandogan, and Zandi (2015) does, in fact, count loops. for which we can adopt a scheme similar to that recently introduced by us Marcato et al. (2023, 2024) to model interacting linear polymer solutions and largely based on the well-known de Gennes’ mapping of the $n\to 0$”-limit of the spin $n$-vector model de Gennes (1979) (the complete derivation of $\Omega_{N}^{{\mathbf{x}}_{0}}$ is given in Sec. S2 in SM). The main idea consists in writing the grand canonical partition function of the system $\Xi$ in the following field-theoretic form,
@@
where $\bm{\varphi}$ and $\bar{\bm{\varphi}}$ are $n$-component field vectors and ${\mathcal{A}}({\bm{\varphi}},\bar{\bm{\varphi}})$ is an effective action (Eq. (S18) in SM, whose validity can be further assessed by applying a version of the famous Wick’s theorem Zinn-Justin (2021) to it (see Sec. S2.1.4 in SM)). While the final expression remains mathematically challenging, it is particularly amenable to an expansion around the saddle-point Orland, Itzykson, and de Dominicis (1985); Doniach, Garel, and Orland (1996); Higuchi (1998); Marcato et al. (2023, 2024), which is obtained by minimizing the effective action. The first term of this expansion corresponds to our MF estimate. Finally, a standard Legendre transform allows us to switch back to the canonical ensemble and get finally $\Omega_{N}^{{\mathbf{x}}_{0}}$. By neglecting all terms up to ${\mathcal{O}}(V^{-1})$, the MF expression for $s=s_{\rho}(\left\{\phi_{f}\right\})$ (Eq. (3)) for finite node density $\rho\equiv N/V$ and as a function of branch-node fractions $\left\{\phi_{f}\right\}\equiv(\phi_{3},...,\phi_{q})$ with $\phi_{f}\equiv N_{f}/N$ is given by
@@
with $e=2.718$... being the Euler number, ()!(...)! the factorial symbol and with the constraints $\phi_{1}=\sum_{f=3}^{q}(f-2)\phi_{f}$ and $\phi_{2}=1-\sum_{f=3}^{q}(f-1)\phi_{f}$ following, respectively, from Eqs. (1) and (2).
@@
i.e. the entropy per node $s_{\rho}/\rho$ is constant. Interestingly, the argument of the logarithm plays now the role of an effective coordination number Lipson and Whittington (1983) which depends on the set of branching probabilities $\{\phi_{f}\}$. In particular, this effective coordination number can take smaller values with respect to the one for linear chains (which, within our MF approximation, is =q=q, i.e. the same as for random walks on the lattice): for instance, by setting $\phi_{3\leq f<q}=0$ and $\phi_{q}\neq 0$, for $q\geq 6$, there are values of $\phi_{q}$ giving an effective coordination number <q<q. Third, of even greater interest is the opposite limit of spanning trees, $\rho=1$, namely at full lattice occupation. In particular, through the Legendre transform of Eq. (5) (see Sec. S4 in SM) it is possible to enumerate the total number of spanning trees on the lattice with unrestricted $N_{f}$. Our MF estimate reproduces the known Wu (1977); Shrock and Wu (2000) exact entropies of spanning trees for different lattices with good accuracy (deviations are no larger than $\simeq 20\%$, see Table S1 in SM). While our approximation is less precise than the Bethe approximation of De Los Rios et al. De Los Rios, Lise, and Pelizzola (2001) (see Table S1 in SM), our method is conceptually superior as it allows us to seize control of the mean number of branch-nodes, a feature absent in De Los Rios, Lise, and Pelizzola (2001).
@@
Figure 2: Distribution of equilibrium densities $\langle\phi_{f}\rangle$ for nodes of functionality $f$ in spanning trees on the triangular lattice ($q=6$). Orange columns correspond to our MF result for an infinite lattice, whereas blue columns refer to the results obtained through the numerical procedure of Pozrikidis Pozrikidis (2016) for a periodic lattice with a total number of 256256 sites.
@@
Figure 3: Asymptotic mean fraction of branch-nodes, $\langle\phi_{3}\rangle$, as a function of the branch chemical potential $\beta\mu_{3}$, comparison between MF prediction Eq. (9) (dashed line) and MC numerical simulations (symbols, corresponding to the values for the largest NN’s reported in Fig. S4 in SM) for different node densities $\rho$ (see colorcode/symbol in the legend) and in spatial dimensions $d=2,3,4$. (Insets) Corresponding ratios between MC results and the MF prediction.
@@
where $\beta=1/(k_{B}T)$. Inverting Eq. (8) for given $\{\mu_{f}\}$, the unknown mean values $\{\langle\phi_{f}\rangle\}$ obey a system of q2q-2 equations which, in general, needs to be solved numerically. However, a general feature easily deduced from Eqs. (5) and (8) is that the values of $\{\langle\phi_{f}\rangle\}$ do not depend on $\rho$. We study first the case of spanning trees (i.e. $\rho=1$) with unrestricted functionality on the triangular lattice (i.e. $q=6$ and $\beta\mu_{f}=0$ for $3\leq f\leq 6$) for which accurate numerical predictions are available Pozrikidis (2016). As shown in Fig. 2, our MF theory agrees very well with the numerical results. Then we consider the situation where $\beta\mu_{f}\neq 0$ and assume that branch-nodes are at most $3$-functional regardless of lattice coordination $q$. With that, Eq. (8) reduces to the following simple expression for $\langle\phi_{3}\rangle$ as a function of the chemical potential $\mu_{3}$:
@@
for tree conformations with $N_{3}$ branch-nodes, weighed through the chemical potential $\beta\mu_{3}$. For node density $\rho<1$, we use the simple and efficient Amoeba algorithm Seitz and Klein (1981); Rosa and Everaers (2016a, b); van der Hoek, Rosa, and Everaers (2024): one of the nodes with $f=1$ is picked at random, removed from the tree and reattached to either one of the “f=1f=1”- or “f=2f=2”-nodes (to preserve the global condition $f\leq 3$), provided that the self-avoiding condition is satisfied. At full lattice occupation ($\rho=1$), where Amoeba is clearly unsuitable, classical algorithms for spanning trees Broder (1989); Aldous (1990) also fail in the present context because they do not allow to weigh conformations according to Eq. (10). To solve this issue, we propose the following new method (for more details, see Sec. S5 in SM): starting from an initial spanning tree, a “f=1f=1”-node is randomly picked and relinked to one of its neighbors provided, again, that the global condition $f\leq 3$ remains satisfied. One can easily see that this algorithm never creates loops in the tree. Trees are then sampled according to (10) by implementing detailed balance through the classical Metropolis rule Metropolis et al. (1953), for values $\beta\mu_{3}=-1,0,1,2$. For each node density $\rho$, we verify that we have effectively reached the asymptotic large-$N$ regime by monitoring the mean branch-node fraction $\langle\phi_{3}\rangle$ as a function of $N$ (see Fig. S4 in SM, notice also the exactness of Eq. (9) for ideal trees (black symbols)). Plots of MF Eq. (9) (lines) and asymptotic MC results (symbols) as a function of $\beta\mu_{3}$, for node density $\rho$ and spatial dimension $d$ are shown in Fig. 3, with the insets reporting the ratios between MC and MF values. We notice that the agreement improves as a function of $\beta\mu_{3}$, $\rho$ and $d$, increasing from a good $\approx 78\%$ in $d=2$, $\rho=0$, $\beta\mu_{3}=-1$ to an excellent $\approx 95\%$ in $d=4$, $\rho=1$, $\beta\mu_{3}=2$. [^3] These findings are in agreement with the known fact Georges and Yedidia (1991) that MF solutions of lattice models become more and more accurate with the increasing of space dimensionality. Moreover, by increasing $\rho$ we expect de Gennes (1979) partial screening of excluded-volume effects and, therefore, a better agreement between simulations and the MF solution, as also seen clearly in our data.
@@
Conclusions – In this work, we have presented a field-theoretic method to study the thermodynamics of single, self-avoiding trees on a generic, regular lattice. We have obtained a compact mean-field expression (Eq. (5)) for the SAT entropy that is, at the same time, significantly simpler and more general than others obtained in seemingly MF contexts Lubensky and Isaacson (1979, 1981); De Los Rios, Lise, and Pelizzola (2001); Zilman and Safran (2002); Wagner, Erdemci-Tandogan, and Zandi (2015), and whose accuracy (see Table S1 in SM) can in principle Bawendi and Freed (1988) be improved by the systematic inclusion of higher order terms beyond the saddle-point approximation utilized here. Notably, our result provides a simple, analytical formula that is able to interpolate between dilute trees ($\rho\ll 1$) and spanning trees ($\rho=1$), as well as between linear and highly branched structures. The results for all these different scenarios have been compared with exact results from the literature, highlighting the good accuracy of our MF approximation. We have also provided predictions for the branching probability (Eq. (9)) that are in good agreement with computational results from Monte Carlo simulations. In this regard, a new simple, yet effective algorithm for the simulation of spanning trees has also been presented. Finally, with little additional effort the model can be extended to include the contribution of nearest-neighbor attractive interactions between tree nodes to model the situation of branched polymers in poor-solvent conditions Rubinstein and Colby (2003). In particular (see Sec. S6 in SM), poor-solvent interactions do not affect the mean values of branch-node fractions found for simple excluded-volume conditions; at the same time, the expected globule-coil transition as a function of temperature is of the same kind of the one for simple linear (i.e. unbranched) polymers in agreement with works Nemirovsky, Dudowicz, and Freed (1992); De Los Rios, Lise, and Pelizzola (2001).
@@
Entropy of self-avoiding branching polymers: Mean-field theory and Monte Carlo simulations
@@
S1 Relation between entropies of undirected and rooted-directed trees
In this Section, we discuss the quantitative relation between the entropies of undirected and rooted-directed SAT’s at the basis of Eq. (3) in main text. For the sake of simplicity, we work with examples relative to the $2d$ square lattice; it is then quite straightforward to extend the arguments to more general lattices.
[…]
Figure S1: Tree conformations (left column) and their corresponding lattice configurations as undirected trees (central column) and rooted-directed trees (right column) on the $3\times 3$ square lattice.
To fix the ideas, consider the tree conformations shown in Fig. S1.
  • (Top panels) The tree conformation does not display rotational symmetry: as a consequence, there are $4$ different lattice configurations of undirected trees corresponding to the same conformation. Then, each of the 4 configurations in the middle column can be translated to any of the sites of the lattice (which is possible because periodic boundary conditions are assumed). Therefore, there are in total $4\times 9=36$ undirected tree lattice configurations corresponding to the same polymer conformation. For what concerns rooted-directed trees, the spatial position of the root has been arbitrarily assigned to the central lattice site. It is then clear that, at fixed root position, the degeneracy of lattice configurations of rooted-directed trees due to rotations is still equal to $4$. However, since we have fixed the position of the root, we have actually reduced the degeneracy due to translations: in fact, the tree can be translated as long as there is always a monomer at the site of the root. Hence, each of the $4$ configurations depicted in the right column has a further translational degeneracy equal to the number of monomers: hence, the total degeneracy in this case is equal to $4\times 5=20$.
  • (Middle panels) Let us now consider the second polymer conformation, which displays partial rotational symmetry. In the undirected case, the degeneracy of lattice configurations due to rotation is now $2$: the degeneracy due to translation remains instead equal to the number of lattice sites, i.e. $9$ in our example. Again, for the rooted-directed lattice configurations the rotational degeneracy is the same as for the undirected case, while the translational degeneracy is equal to the number of monomers, i.e. $3$.
  • (Bottom panels) The last conformation is fully symmetric under rotations on the square lattice, therefore there is no degeneracy due to rotations for the lattice configurations, neither in the undirected nor in the rooted-directed case. However, the degeneracy due to translations remains equal to $9$ for the undirected case and equal to $5$ for the rooted-directed case (i.e. the number of sites and the number of monomers, respectively).
By these examples, it should be clear that the degeneracy of each polymer conformation is proportional to $V$ for the undirected case and to $N$ for the rooted-directed case, with the proportionality constant depending on the symmetry of the conformations: however, at fixed position of the root, the proportionality constant is the same in both cases. In other words, we have that
$\frac{\Omega_{N}}{\Omega_{N}^{{\mathbf{x}}_{0}}}=\frac{V}{N}\,.$ (S1)
By taking the logarithm of both sides of Eq. (S1) and dividing by $V$, Eq. (3) in main text follows easily.
S2 Field-theoretic partition function and mean-field approximation
In this Section, we illustrate in detail the derivation of the mean-field expression of the entropy (Eq. (5) in main text) for self-avoiding trees (SAT’s).
S2.1 Spin $n$-vector model for SAT’s and $n\to 0$ limit
S2.1.1 Formulation of the problem
We consider a regular lattice with coordination number (equal to the number of nearest neighbors per lattice site) =q=q, total number of sites =V=V and with periodic boundary conditions; lengths are expressed in units of the lattice step length. As explained in main text, the goal is to estimate $\Omega_{N}^{{\mathbf{x}}_{0}}(\left\{N_{f}\right\})$, the total number of rooted-directed trees with root on the lattice site ${\mathbf{x}}_{0}$, $N$ nodes and $\left\{N_{f}\right\}\equiv(N_{3},...,N_{q})$ branch-nodes of functionality $f$.